# Unified Triplet-Level Hallucination Evaluation for Large Vision-Language Models

### v2

@@ Abstract
Despite the outstanding performance in vision-language reasoning, Large Vision-Language Models (LVLMs) might generate hallucinated contents that do not exist in the given image. Most existing LVLM hallucination benchmarks are constrained to evaluate the object-related hallucinations. However, the potential hallucination on the relations between two objects, i.e., relation hallucination, still lacks investigation. To remedy that, in this paper we design a unified framework to measure object and relation hallucination in LVLMs simultaneously. The core idea of our framework is to conduct hallucination evaluation on (object, relation, object) triplets extracted from LVLMs’ responses, and thus, could be easily generalized to different vision-language tasks. Based on our framework, we further introduce Tri-HE, a novel Triplet-level Hallucination Evaluation benchmark which can be used to study both object and relation hallucination at the same time. We conduct comprehensive evaluations on Tri-HE and observe that the relation hallucination issue is even more serious than object hallucination among existing LVLMs, highlighting a previously neglected problem towards reliable LVLMs. Moreover, based on our findings, we design a simple yet effective training-free approach to mitigate hallucinations for LVLMs, with which, we exceed all open-sourced counterparts on Tri-HE, achieving comparable performance with the powerful GPT-4V. Our dataset and code for the reproduction of our experiments are available publicly at https://github.com/wujunjie1998/Tri-HE.
@@ 1 Introduction
Large Vision-Language Models (LVLMs) [1, 2, 3, 4] have attracted significant attention. Despite the superior performances, existing works primarily focus on enhancing the helpfulness of LVLMs without careful consideration of the reliability of responses generated by LVLMs. However, it has already been observed by recent literature that LVLMs suffer from severe hallucination [5, 6, 7, 8, 9], i.e., LVLMs might generate contents that do not exist in the given image, probably due to insufficient training during visual instruction tuning. A typical example is provided in Figure 1(a), where the LLaVA [2] model considers the location to be busy, simply because LLaVA recognizes that it is a train station with several people existing.
With the prevalence of LVLMs, enormous works have started to explore the evaluation and analysis of LVLM hallucination. However, two problems are observed: 1) Hallucination category: most existing works focus on object-related hallucination [5, 6, 10] (i.e., LVLMs describing an object not existing in the given image) while ignoring the possibility that even when two objects are successfully recognized, LVLMs might still mess up with their relationships when conducting commonsense reasoning. As illustrated in the example in Figure 1(a), LLaVA successfully recognizes the “people” and the train station “area”, yet predicts their relation to be “walking around” that cannot be directly obtained from the given image. Therefore, a unified definition and taxonomy is necessary to integrate different kinds of LVLM hallucination.
2) Hallucination discrimination: existing works [5, 6, 8, 11] rely on self-discrimination (e.g., Yes/No questions) to testify LVLMs’ awareness of objects and relations in the given image. When prompted with an instruction “Is there a {content} in the image?”, hallucination is determinied only when LVLMs answer “Yes.” for a non-existing content. However, these benchmarks require transforming general vision-language tasks into "yes/no" formats, limiting their applicability. Additionally, different LVLMs may have different capability in answering such“yes/no” questions since they are pre-trained on different data, which may make the evaluation results biased [12]. So, here arises our question: can we perform a unified and unbiased evaluation for various kinds of hallucination?
To this end, this paper first proposes a unified framework to measure both object and relation hallucinations (§3). Specifically, it first extracts knowledge graphs as triplets from LVLMs’ responses, then uses external judges to compare these triplets with scene graphs of input images for hallucination evaluation. In this way, our method can provide fine-grained evaluation results on the response to any vison-language tasks. Based on the proposed framework, we then construct Tri-HE, a novel Triplet-level Hallucination Evaluation benchmark to evaluate object and relation hallucination simultaneously (§4). As reported in §5 and Figure 1(b), relation hallucination is indeed a common nightmare even for the powerful GPT-4V model [13], and is even more severe than object hallucination. By comparing the performance of different LVLMs, we further recognize several findings that could potentially decrease LVLM hallucination (§5.2). In addition, our triplet-level hallucination judge empowered by GPT-4 reveals remarkable alignment with human judgments (Table 3). This motivates us to integrate triplet description into LVLMs and propose a simple yet effective training-free hallucination mitigation method (§5.4). With this method, the LLaVA-1.5 model [14] exceeds all the open-sourced counterparts, achieving comparable performance with the powerful GPT-4V, as shown in Figure 1(b) and Table 5.
Our main contributions contain three parts:
  1. 1.
    We introduce a unified framework to measure both the object and relation hallucination of LVLMs simultaneously. Specifically, we propose to conduct a triplet-level hallucination evaluation, which can provide more accurate and fine-grained evaluation results.
  2. 2.
    Based on the framework, we propose Tri-HE, a novel triplet-level fine-grained hallucination evaluation benchmark for LVLMs.
  3. 3.
    We propose a simple yet effective training-free hallucination mitigation method that exceeds all open-sourced counterparts, performing on par with GPT-4V on Tri-HE.
@@
Figure 1: Overview of Tri-HE evaluation pipeline. (a) With the provision of images, scene graphs, and questions, knowledge graphs (i.e., triplets) are extracted from LVLM responses, which are then judged by GPT-4. (b) The radar plot showcases the evaluation results among different LVLMs (lower values demonstrate fewer hallucinations).
@@ 2.1 Large Vision-Language Models (LVLMs)
The powerful capability exhibited by Large Language Models (LLMs) has facilitated the extension of LLMs towards the multi-modal domain. LLMs are empowered to understand and reason about both images and text by aligning representations from visual encoders to pre-trained language models, followed by visual instruction tuning. LLaVA [14, 2] proposes to use a simple projection layer to integrate the visual representations into textual encoders, which is further enhanced in Shikra [15] by incorporating referential dialogue tasks. Instead, BLIP [16] proposes the Q-Former architecture to extract useful information from the visual representations, which is also used by MiniGPT-4 [3] and InstructBLIP [1]. MoCLE [17] further introduces the Mixture-of-Experts architecture into LVLMs to deal with the data conflict during instruction tuning. Although powerful, existing works primarily focus on improving the helpfulness, without a thorough analysis on the reliability of LVLMs.
@@ 2.2 Hallucination Evaluation in LVLMs
With the prevalence of LVLMs, a growing number of studies have been conducted on their hallucination issues [9, 18, 19, 20, 21, 6, 8, 22]. Previous hallucination evaluation works can be categorized into two groups: 1) solely evaluating object hallucinations or do not distinguish different hallucinations [23, 24, 6, 10], which neglects other hallucination types like relation hallucination and is thus not comprehensive. The other type of works use “yes/no” questions to evaluate LVLM’s relation/object hallucinations [5, 25, 8, 11]. However, these benchmarks require transforming general vision-language tasks into “yes/no” formats, limiting their applicability. Also, different LVLMs may have different ability in answering such“yes/no” questions since they are pre-trained on different data, which may bias the evaluation results. To remedy this research gap, our paper proposes a triplet-level evaluation framework that can provide fine-grained object and relation hallucinations for responses to any vision-language tasks, with an evaluation benchmark Tri-HE that incorporates questions requiring more complicated commonsense reasoning.
It is worth noting that a concurrent benchmark, Reefknot [26], also evaluates relation hallucinations at the triplet level. However, compared to Tri-HE, Reefknot has several limitations: 1) It fully relies on existing datasets, which may lead to data contamination, whereas Tri-HE consists of questions synthesized by GPT-4V. 2) Reefknot evaluates only a limited number of relation types, while Tri-HE covers a broader range, enabling more comprehensive evaluation results as shown in §5. 3) Reefknot employs only one entailment-based hallucination discriminator, whereas Tri-HE offers more accurate hallucination detection using GPT-4, which can simultaneously identify both object and relation hallucinations.
3 Tri-HE Framework Formulation
Inspired by the relation extraction [27] tasks in NLP, in this section, we propose a unified framework to evaluate both object and relation hallucinations via the object-relation triplets (i.e., $(\text{Object}_{1},\text{Relation},$ $\text{Object}_{2})$). Here the objects and relations can either be a word or a phrase with attributes. We start by defining object and relation hallucinations via triplets in §3.1, based on which, we define our evaluation metrics and pipeline in §3.2 and §3.3 separately.
@@ 3.1 Definitions
As shown in Figure 1(a), given an input image $I$, a question $Q$ related to $I$ with its ground truth answer $A$, and the answer $A_{\theta}$ predicted by a LVLM $A_{\theta}(\cdot|Q,I)$ parameterized with $\theta$, we first define,
@@
  • $G^{\prime}=(V^{\prime}\subseteq V,E^{\prime}\subseteq E)$ as the knowledge graph that includes all the required objects and relations to answer $Q$.
@@
Note that here all graphs can be converted to a set of triplets (i.e., $G=\{(v_{1},e,v_{2})\}$, where $v_{1},v_{2}\in V$ and $e\in E$). A common nightmare in previous LVLM hallucination literature lies in the ambiguous discrimination between prediction hallucinations and errors [28]. To obtain unbiased hallucination evaluation results, we separate them depending on whether or not the wrongly generated objects or relations exist in the given image $I$. Specifically, given a triplet $(v_{1},e,v_{2})\in G_{\theta}$, we have the following definitions,
  • Object hallucination: if $v_{1}\notin V$ or $v_{2}\notin V$, suggesting $A_{\theta}$ includes an object not within $I$. For example, the triplet (location, suggests, popular spot for socializing) in Figure 1(a) encounter an object hallucination since the object “ popular spot for socializing” cannot be obtained from $V$.
@@ 3.2 Evaluation Metrics
With the above definition in hand, given the knowledge graph $G_{\theta}$ extracted from a model response $A_{\theta}$, we calculate hallucination rates of $A_{\theta}$ as the proportion of hallucinated triplets in $G_{\theta}$. Most previous works (e.g., POPE [5]) directly evaluate the hallucination rate at the object-level with respect to the total number of predicted objects, yet make their results not comparable among LVLMs, since different LVLMs might refer to different numbers of objects in their responses. To address this issue, we instead opt to calculate the hallucination rate in the question- and image-level. Specifically, we calculate two types of hallucination rates, including the question-level hallucination rate ($\text{Hallu}_{\text{Q}}$) and image-level hallucination rate ($\text{Hallu}_{\text{I}}$), as defined in the following,
@@ 3.3 Evaluation Pipeline
With the definitions and evaluation metrics provided in §3.1 and §3.2, the remaining problems contain two parts: 1) how to extract the knowledge graph $G_{\theta}$ from LVLM responses $A_{\theta}$, and 2) how to judge a triplet in $G_{\theta}$ is hallucinated or not. The overview of our pipeline is illustrated in Figure 1(a).
@@
Given an LVLM response $A_{\theta}$ with the corresponding question $Q$ and image $I$, we extract the knowledge graph $G_{\theta}$ from $A_{\theta}$ via prompting GPT-4. Check our prompt for knowledge graph extraction in Figure 6. Afterward, we propose two different strategies to judge whether a triplet $(v_{1},e,v_{2})\in G_{\theta}$ includes hallucination based on the ground truth answer $A$ and the image scene graph $G$, as described in the following.

### v3

@@ Abstract
Despite the outstanding performance in vision-language reasoning, Large Vision-Language Models (LVLMs) might generate hallucinated contents that do not exist in the given image. Most existing LVLM hallucination benchmarks are constrained to evaluate the object-related hallucinations. However, the potential hallucination on the relations between two objects, i.e., relation hallucination, still lacks investigation. To remedy that, we design a unified framework to measure the object and relation hallucination in LVLMs simultaneously. The core idea of our framework is to evaluate hallucinations via (object, relation, object) triplets extracted from LVLMs’ responses, making it easily generalizable to different vision-language tasks. Based on our framework, we further introduce Tri-HE, a novel Triplet-level Hallucination Evaluation benchmark which can be used to study both object and relation hallucination at the same time. With comprehensive evaluations on Tri-HE, we observe that the relation hallucination issue is even more serious than object hallucination among existing LVLMs, highlighting a previously neglected problem towards reliable LVLMs. Moreover, based on our findings, we design a simple training-free approach that effectively mitigates hallucinations for LVLMs. Our dataset and code for the reproduction of our experiments are available publicly at https://github.com/wujunjie1998/Tri-HE.
@@ 1 Introduction
Large Vision-Language Models (LVLMs) (Dai et al., 2023; Liu et al., 2023b; Chen et al., 2024a; Cai et al., 2024) have attracted significant attention. Despite the superior performances, existing works primarily focus on enhancing the helpfulness of LVLMs without careful consideration of the reliability of responses generated by LVLMs. However, it has already been observed by recent literature that LVLMs suffer from severe hallucination (Li et al., 2023e; Wang et al., 2023b; c; Guan et al., 2024; Chen et al., 2024b), i.e., LVLMs might generate contents that do not exist in the given image, probably due to insufficient training during visual instruction tuning. A typical example is provided in Figure 1(a), where the LLaVA (Liu et al., 2023b) model considers the location to be busy, simply because LLaVA recognizes that it is a train station with several people existing but without reasoning about their relationships.
@@
Figure 1: Overview of the unified hallucination evaluation pipeline of Tri-HE. (a) With the provision of images, scene graphs, and questions, knowledge graphs (i.e., triplets) are extracted from LVLM responses, which are then judged by an LLM (GPT-4 here). (b) The radar plot showcases the evaluation results among different LVLMs (lower values demonstrate fewer hallucinations).
With the prevalence of LVLMs, enormous works have started to explore the evaluation and analysis of LVLM hallucination. However, two problems are observed: 1) Hallucination category: most existing works focus on object-related hallucination (Li et al., 2023e; Wang et al., 2023b; Chen et al., 2024c) (i.e., LVLMs describing an object not existing in the given image) while ignoring the possibility that even when two objects are successfully recognized, LVLMs might still mess up with their relationships when conducting commonsense reasoning (Gou et al., 2025a). As illustrated in the example in Figure 1(a), LLaVA successfully recognizes the “people” and the train station “area”, yet predicts their relation to be “walking around” that cannot be directly obtained from the given image. Therefore, a unified definition and taxonomy is necessary to integrate different kinds of LVLM hallucination.
2) Hallucination discrimination: To evaluate how severe LVLMs hallucinate objects and their relationships within given images, prior works generally use either self-discrimination methods (e.g., Yes/No questions) (Li et al., 2023e; Wang et al., 2023b; Guan et al., 2024; Wu et al., 2024b) or template-driven discrimination approaches (e.g., “What is the relation with A and B?”) such as Reefknot (Zheng et al., 2024). However, such methods inherently constrain LVLMs to generate short answers like “Yes/No” or “A has {} relation to B”. Given that LVLMs have varying capabilities to produce brief responses due to differences in pre-training datasets, this could introduce biased evaluation results (Chen et al., 2023b; Liu et al., 2024). For instance, Li et al. (2023e) have shown that InstructBLIP (Dai et al., 2023) tends to produce shorter outputs compared to other LVLMs, thus inflating its performance in answering the above type of questions and leads to hallucination evaluation bias. Moreover, these benchmarks require transforming general vision-language tasks into specific formats like "Yes/No", limiting their applicability. Therefore, we raise the following research question: Can we develop a unified and unbiased evaluation framework capable of evaluating various types of hallucinations in LVLM responses across diverse tasks?
To this end, we first propose a unified framework to simultaneously measure object and relation hallucinations in LVLM responses (§3). Specifically, our framework extracts knowledge graphs represented as triplets from LVLM-generated responses and then employs external evaluators to compare these triplets against the corresponding scene graphs from the input images. Consequently, our method facilitates hallucination evaluation for responses across diverse vision-language tasks, independent of the specific question formats. Leveraging this unified framework, we further introduce the Tri-HE, a novel benchmark for Triplet-level Hallucination Evaluation, designed explicitly to assess both object and relation hallucinations (§4). Our experimental findings presented in §5 and Figure 1(b) confirm that relation hallucination poses a significant challenge for both closed-source and open-source LVLMs, often surpassing object hallucination in severity. By systematically comparing LVLMs’ performance, we identify key insights that could potentially reduce hallucination rates (§5.2). Furthermore, our proposed triplet-level hallucination judge, powered by LLMs, demonstrates impressive alignment with human judgments (Table 3). Motivated by these observations, we incorporate explicit triplet descriptions into LVLM prompts and introduce a straightforward yet effective training-free method to mitigate hallucinations (§5.4).
Our primary contributions are summarized in the following three perspectives:
  1. 1.
    We propose a unified framework capable of jointly evaluating object and relation hallucinations in LVLM responses across diverse vision-language tasks. In particular, our triplet-level evaluation offers a finer-grained, more accurate assessment compared to existing methods.
  2. 2.
    Building upon this framework, we introduce Tri-HE, a novel triplet-level fine-grained hallucination evaluation benchmark tailored specifically for LVLMs.
  3. 3.
    We propose a simple yet highly effective training-free hallucination mitigation approach that surpasses the open-source LVLM competitors.
@@ 2.1 Large Vision-Language Models (LVLMs)
The powerful capability exhibited by Large Language Models (LLMs) has facilitated the extension of LLMs towards the multi-modal domain. LLMs are empowered to understand and reason about both images and text by aligning representations from visual encoders to pre-trained language models, followed by visual instruction tuning. LLaVA (Liu et al., 2023a; b) proposes to use a simple projection layer to integrate the visual representations into textual encoders, which is further enhanced in Shikra (Chen et al., 2023d) by incorporating referential dialogue tasks. Instead, BLIP (Li et al., 2023a) proposes the Q-Former architecture to extract useful information from the visual representations, which is also used by MiniGPT-4 (Zhu et al., 2023) and InstructBLIP (Dai et al., 2023). InternLM (Dong et al., 2024) aligns with more diverse instruction data with the conditional online reinforcement learning from human feedback (RLHF) strategy, while MoCLE (Gou et al., 2023) further introduces the Mixture-of-Experts architecture into LVLMs to deal with the data conflict during instruction tuning. Although powerful, existing works primarily focus on improving the helpfulness and robustness Gou et al. (2025b), without a thorough analysis of the reliability of LVLMs.
@@ 2.2 Hallucination Evaluation in LVLMs
With the prevalence of LVLMs, a growing number of studies have been conducted on their hallucination issues (Chen et al., 2024b; d; Han et al., 2024; Huang et al., 2023; Li et al., 2023b; Wang et al., 2023b; Guan et al., 2024; Yue et al., 2024). Previous hallucination evaluation works can be categorized into two groups: 1) solely evaluating object hallucinations or do not distinguish different hallucinations (Zhao et al., 2023; Li et al., 2023c; Wang et al., 2023b; Chen et al., 2024c), which neglects other hallucination types like relation hallucination and is thus not comprehensive. The other type of works use “yes/no” questions to evaluate LVLM’s relation/object hallucinations (Li et al., 2023e; Wang et al., 2023a; Guan et al., 2024; Wu et al., 2024b). However, these benchmarks require transforming general vision-language tasks into “yes/no” formats, limiting their applicability. Also, different LVLMs may have different ability in answering such“yes/no” questions since they are pre-trained on different data, which may bias the evaluation results. To remedy this research gap, our paper proposes a triplet-level evaluation framework that can provide fine-grained object and relation hallucinations for responses to any vision-language tasks, with an evaluation benchmark Tri-HE that incorporates questions requiring more complicated commonsense reasoning.
It is noteworthy that a concurrent benchmark, Reefknot (Zheng et al., 2024), similarly assesses relation hallucinations at the triplet level. However, Reefknot exhibits several limitations compared to Tri-HE. First, Reefknot constructs VQA questions based on a simple template, “What is the relation between A and B?”, restricting both the variety of vision-language tasks that can be evaluated and the length of LVLM-generated responses, potentially introducing evaluation biases, similarly with “yes/no” questions for evaluating object hallucination. In contrast, our framework is flexible enough to be applied to various vision-language tasks. Moreover, since the questions in Tri-HE are generated by GPT-4V, it can cover a wider range of relation types compared to template-based questions, thus providing more comprehensive evaluation results. Second, Reefknot relies solely on a single entailment-based hallucination discriminator, whereas Tri-HE leverages powerful LLM-based discriminators capable of accurately and simultaneously identifying both object and relation hallucinations, leading to more comprehensive hallucination evaluation results.
3 Unified Hallucination Evaluation Framework Formulation
Inspired by the relation extraction (Xiaoyan et al., 2023) task, in this section, we propose a unified framework to evaluate both object and relation hallucinations via the object-relation triplets (i.e., $(\text{Object}_{1},\text{Relation},$ $\text{Object}_{2})$). Here the objects and relations can either be a word or a phrase with attributes. We start by defining object and relation hallucinations via triplets in §3.1, based on which, we define our evaluation metrics and pipeline in §3.2 and §3.3 separately.
@@ 3.1 Definitions
As illustrated in Figure 1(a), we formulate our framework with the standard VQA setting (although they can be generalized to evaluate hallucinations in any vision-language tasks given available scene graph annotations, as discussed in §3.4). Specifically, considering an input image $I$, a corresponding question $Q$ associated with image $I$, its ground truth answer $A$, and the answer $A_{\theta}$ predicted by an LVLM $A_{\theta}(\cdot|Q,I)$ parameterized by $\theta$, we can first have the following definitions:
@@
  • $G^{\prime}=(V^{\prime}\subseteq V,E^{\prime}\subseteq E)$ as the knowledge graph that includes all the required objects and relations to answer $Q$ precisely.
@@
Note that here all graphs can be converted to a set of triplets (i.e., $G=\{(v_{1},e,v_{2})\}$, where $v_{1},v_{2}\in V$ and $e\in E$). A common nightmare in previous LVLM hallucination literature lies in the ambiguous discrimination between prediction hallucinations and errors (Ji et al., 2023). To obtain unbiased hallucination evaluation results, we separate them depending on whether or not the wrongly generated objects or relations exist in the given image $I$. Specifically, given a triplet $(v_{1},e,v_{2})\in G_{\theta}$, we have the following definitions,
  • Object hallucination: if $v_{1}\notin V$ or $v_{2}\notin V$, suggesting $A_{\theta}$ includes an object not within $I$. For example, the triplet (location, suggests, popular spot for socializing) in Figure 1(a) encounters an object hallucination since the object “ popular spot for socializing” cannot be obtained from $V$.
@@ 3.2 Evaluation Metrics
With the above definition in hand, given the knowledge graph $G_{\theta}$ extracted from a model response $A_{\theta}$, we calculate the hallucination rates of $A_{\theta}$ as the proportion of hallucinated triplets in $G_{\theta}$. Most previous works (e.g., POPE (Li et al., 2023e)) directly evaluate the hallucination rate at the object-level with respect to the total number of predicted objects, yet make their results not comparable among LVLMs, since different LVLMs might refer to different numbers of objects in their responses. To address this issue, we instead opt to calculate the hallucination rate at the question- and image-level. Specifically, we calculate two types of hallucination rates, including the question-level hallucination rate ($\text{Hallu}_{\text{Q}}$) and image-level hallucination rate ($\text{Hallu}_{\text{I}}$), as defined in the following,
@@ 3.3 Evaluation Pipeline
With the definitions and evaluation metrics provided in §3.1 and §3.2, the remaining problems contain two folds: 1) how to extract the knowledge graph $G_{\theta}$ from LVLM responses $A_{\theta}$, and 2) how to judge whether a triplet in $G_{\theta}$ is hallucinated or not. The overview of our pipeline is illustrated in Figure 1(a).
@@
Given an LVLM response $A_{\theta}$ with the corresponding question $Q$ and image $I$, we extract the knowledge graph $G_{\theta}$ from $A_{\theta}$ via prompting GPT-4. Check our prompt for knowledge graph extraction in Appendix §A.1. Afterwards, we propose two different strategies to judge whether a triplet $(v_{1},e,v_{2})\in G_{\theta}$ includes hallucination based on the ground truth answer $A$ and the image scene graph $G$, as described in the following.
NLI Judge.
The first strategy is implemented with a natural language inference (NLI) (Reimers & Gurevych, 2019) model  [^1] https://huggingface.co/sentence-transformers/all-mpnet-base-v2. Specifically, given an extracted triplet, we first calculate its cosine similarity scores with all triplets in the image scene graph $G$ and only retain those ground truth (GT) triplets with similarity scores greater than 0.5 to refine the information that will be used for the NLI model. If no triplets in $G$ meet this criterion, only the top three GT triplets with the highest similarity scores will be kept, which are then taken as ground truth inputs for the NLI model to make predictions. If the NLI score between the extracted triplet and ground truth triplets is lower than 0.6, suggesting the extracted triplet cannot be induced based on GT triplets, and therefore, resulting in a hallucination.
To determine the threshold, we randomly selected question instances from 10 images and reviewed the set of filtered triplets that were returned. The similarity score threshold was adjusted to 0.5 for the most reasonable returned triplets. These triplets later concatenate together as the ground truth required for generating NLI judgments. In determining if a generated triplet was hallucinated, we further review the NLI judgment results in different thresholds, ultimately deciding on a threshold of 0.6.
LLM Judge.
Another evaluation strategy is to leverage prompting of a powerful LLM, a widely-adopted practice in recent works for assessing LLM outputs (Zheng et al., 2023). In this work, we primarily utilize GPT-4 in LLM judge to determine whether a given extracted triplet $(v_{1},e,v_{2})\in G_{\theta}$ can be directly obtained or inferred from the image scene graph $G$. Note that:
  1. 1.
    We do not employ GPT-4V in LLM judge, as Li et al. (2024) have reported that the text-only GPT-4 judge is more consistent with human preferences than GPT-4V judge.
  2. 2.
    Frontier open-source models, such as LLaMA-3.3, can similarly deliver reliable and cost-efficient hallucination evaluation results (see detailed analysis in Table 3), which might be superior alternatives when the GPT-4 judge is unavailable.
Additionally, if a triplet $(v_{1},e,v_{2})$ is judged as hallucinated, we further prompt the LLM judge to clarify whether the hallucination pertains specifically to the relation $e$ or the objects $v_{1},v_{2}$. Refer to Appendix §A.2 for the prompt of LLM judge in our experiments  [^2] For both knowledge graph extraction and LLM judge, we utilize the “gpt-4-1106-preview” model via OpenAI’s API with default inference parameters..
3.4 Generalizability of our Framework
Although we formulate our unified triplet-level hallucination evaluation framework in the sections mentioned above primarily based on the VQA tasks, it is capable of evaluating hallucinations in LVLM responses for any vision-language task (since knowledge graph extraction is suitable for all natural-language-based responses), provided that the corresponding scene graphs for the test images are available or extracted by pre-trained expert models. This underscores the task-agnostic design of our proposed framework and highlights its strong generalization capability. We leave the detailed exploration for other vision-language tasks for future work.
4 Tri-HE Construction
Following the formulation in §3, in this section, we provide a detailed discussion on how to construct our benchmark Tri-HE for a unified triplet-level evaluation of both hallucinations in LVLMs.
Image Collection.
The construction of Tri-HE begins with images from the GQA dataset (Hudson & Manning, 2019), as the scene graph annotations provided by GQA naturally fit our triplet-level hallucination evaluation formulation. Nevertheless, some scene graphs in GQA contain incomplete object relationships, omitting information necessary for accurate question answering. To mitigate this issue, we initially filter the GQA images, retaining only those whose scene graphs contain at least five object relations (edges between nodes). Subsequently, we select 300 images from the filtered images according to the following criteria:
  1. 1.
    Each image must contain more than two related objects.
  2. 2.
    Each image must be sufficiently clear to discern all visual details.
This procedure ensures a set of high-quality images suitable for subsequent dataset construction.
VQA Question Generation
Next, since the VQA questions in the GQA dataset have already been extensively used during the pre-training of current LVLMs, we instead employ GPT-4V  [^3] We use the “gpt-4-vision-preview” model here, the same as in §5.2. to generate novel question-answer pairs for each image to avoid data contamination. To effectively examine both object and relation hallucinations in LVLM responses, we aim to generate questions that necessitate commonsense reasoning grounded on the provided images. Specifically, for every image, GPT-4V is prompted to generate 10 questions along with their answers  [^4] Check §A.3 for the prompt used here., each requiring image-based commonsense reasoning to be answered. Furthermore, we ask GPT-4V to produce relation triplets describing the reasoning processes of answering the questions. These additional triplets can subsequently be used to enrich the original image scene graphs.
VQA Question Verification
Following the initial generation of VQA questions, three annotators manually examine all generated questions, answers, and triplets based on the following criteria:
  1. 1.
    Each question must be valid and answerable using commonsense reasoning based on the input image.
  2. 2.
    Each answer must appropriately address the question using commonsense reasoning.
  3. 3.
    Each triplet must accurately describe the corresponding answer and must only contain objects visible within the image.
Questions or answers failing to meet these conditions are discarded, while invalid triplets are excluded from the respective scene graphs. To validate the annotation consistency, the annotators jointly annotate 100 common question-answer pairs, achieving a Krippendorff’s alpha (Krippendorff, 2011) of 0.62, demonstrating substantial inter-annotator agreement and guaranteeing the consistency of Tri-HE annotations.
Statistics.
The overall statistics for Tri-HE are summarized in Table 1. As described in Figure 2, each image in Tri-HE is linked to a scene graph and a set of question-answer pairs that require reasoning, accompanied by ground truth triplet annotations. Note that since the quality of each question in Tri-HE is manually verified, expanding its size requires significant resources and poses challenges. Nonetheless, the number of images and questions in Tri-HE is comparable to existing LVLM hallucination evaluation benchmarks such as Zhao et al. (2023) and Guan et al. (2024). Furthermore, as demonstrated in Table 3, Tri-HE has been sufficient to provide reliable hallucination evaluation results, largely thanks to the high-quality annotation procedure described above.
#Images #Questions #Objects #Relations #Questions/Image #Triplets/SG
300 1226 1723 618 4.09 19.10
Table 1: Statistics of Tri-HE. “SG” refers to Scene Graph.